OpenStudy (anonymous):

1. Is the sequence geometric? If so, identify the common ratio. 6, 12, 24, 48, ... (1 point)

OpenStudy (anonymous):

http://www.weegy.com/?ConversationId=59B6EB0C

OpenStudy (anonymous):

That should answer your question.

OpenStudy (anonymous):

Could you answer mine?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

I just bumped it to the top. :)

OpenStudy (anonymous):

yes;2 6*2=12 12*2=24 24*2=48